# On the Curses of Future and History in Future-dependent Value Functions for Off-policy Evaluation

### v1

@@ 3 Future-dependent Value Functions
$|J(\pi_{e})-$ $\mathbb{E}_{\mathcal{D}}[\widehat{V}(f_{1})]|\leq cH\max\{C_{\mathcal{V}}+1,C_{\Xi}\}$
$~{}\cdot\mathrm{IV}(\mathcal{V})\mathrm{Dr}_{\mathcal{V}}[d^{\pi_{e}},d^{\pi_{b}}]\sqrt{\frac{C_{\mu}\log\frac{|\mathcal{V}||\Xi|}{\delta}}{n}},$
where $c$ is an absolute constant, [^3] The value of $c$ can differ in each occurrence, and we reserve the symbol $c$ for such absolute constants. and
$\mathrm{IV}(\mathcal{V})$ $:=\max_{h}\sup_{V\in\mathcal{V}}\sqrt{\frac{\mathbb{E}_{\pi_{b}}\left[\left(\mathcal{B}^{{\mathcal{S}}}V\right)(s_{h})^{2}\right]}{\mathbb{E}_{\pi_{b}}\left[\left(\mathcal{B}^{\mathcal{H}}V\right)(\tau_{h})^{2}\right]}},$
$\mathrm{Dr}_{\mathcal{V}}[d^{\pi_{e}},d^{\pi_{b}}]$ $:=\max_{h}\sup_{V\in\mathcal{V}}\sqrt{\frac{\mathbb{E}_{\pi_{e}}\left[\left(\mathcal{B}^{{\mathcal{S}}}V\right)(s_{h})^{2}\right]}{\mathbb{E}_{\pi_{b}}\left[\left(\mathcal{B}^{{\mathcal{S}}}V\right)(s_{h})^{2}\right]}},$
$C_{\mathcal{V}}$ $:=\max_{V\in\mathcal{V}}\|V\|_{\infty},~{}~{}C_{\Xi}:=\max_{\xi\in\Xi}\|\xi\|_{\infty}.$
See proof in Appendix B.2. Among the objects that appear in the bound, some are mundane and expected (e.g., horizon $H$, the complexities of function classes $\log|\mathcal{V}||\Xi|$, etc.). We now focus on the important ones, which reveals the open questions we shall investigate next:
@@
    $\mathrm{Dr}_{\mathcal{V}}[d^{\pi_{e}},d^{\pi_{b}}]$ is the term that measures the coverage of $\pi_{b}$ over $\pi_{e}$ on the latent state space ${\mathcal{S}}$, as the expression inside square-root can be bounded by $\max_{s_{h}}d^{\pi_{e}}(s_{h})/d^{\pi_{b}}(s_{h})$. When $d^{\pi_{b}}$ induces a uniform distribution over ${\mathcal{S}}_{h}$, this is bounded by $S$, which we consider as a benign scaling in our problem.
@@
    (Q2) $\mathrm{IV}(\mathcal{V})$ appears because we use $\mathcal{B}^{\mathcal{H}}V$ as a proxy for $\mathcal{B}^{{\mathcal{S}}}V$, which is only a linear measure of the latter, $(\mathcal{B}^{\mathcal{H}}V)(\tau_{h})=\langle\mathbf{b}(\tau_{h}),\mathcal{B}^{{\mathcal{S}}}_{h}V\rangle$. $\mathrm{IV}(\mathcal{V})$ reflects the conversion ratio between them. Uehara et al. (2022) pointed out that the value is finite if $M_{\mathcal{H},h}$ has full-row rank ($S$), but a quantitative understanding is missing.
In the rest of the paper, we propose new coverage assumptions, analyses, and algorithms to answer the above questions, which provides deeper understanding on the mathematical structure of OPE in POMDPs.
@@ 4 Boundedness of FDVFs
We start with Q1: when are FDVFs bounded? Recall from Definition 3 that a FDVF at level $h$, $V_{\mathcal{F},h}$, is any function satisfying:
$M_{\mathcal{F},h}\times V_{\mathcal{F},h}=V_{{\mathcal{S}},h}^{\pi_{e}},$ (6)
where $M_{\mathcal{F},h}$ is the outcome matrix with $(M_{\mathcal{F},h})_{i,j}=\operatorname{\Pr}_{\pi_{b}}(f_{h}=j\mid s_{h}=i)$. Since the equation can have many solutions and we only need to find one of them, it suffices to provide an explicit construction of a $V_{\mathcal{F},h}$ and show it is well bounded. Also note that the equations for different $h$ are independent of each other, so we can construct $V_{\mathcal{F},h}$ for each $h$ separately.
We first describe two natural constructions & analyses, both of which yield exponentially large upper bounds.
@@
Uehara et al. (2022) provided a cruel bound on $\|V_{\mathcal{F}}\|_{\infty}$ which scales with $1/\sigma_{\min}(M_{\mathcal{F},h})$, which we show shortly is an exponential-in-length quantity. However, it is not hard to notice that in certain benign cases, $V_{\mathcal{F}}$ does not have to blow-up exponentially and has obviously bounded constructions.
As a warm-up, consider the on-policy case of $\pi_{b}=\pi_{e}$, where a natural solution is
$V_{\mathcal{F}}(f_{h})=R(f_{h}):=\sum_{h^{\prime}=h}^{H}R(o_{h^{\prime}},a_{h^{\prime}}).$
Here we use a slight abuse of notation $R(f_{h})$ to simply add up the Monte-Carlo rewards in future $f_{h}$, which is bounded in $[0,H]$. The construction also trivially satisfies $\mathbb{E}_{\pi_{b}}\left[V_{\mathcal{F}}(f_{h})\mid s_{h}\right]=V_{{\mathcal{S}}}^{\pi_{e}}(s_{h})$ when $\pi_{b}=\pi_{e}$. In the more general setting of $\pi_{b}\neq\pi_{e}$, the hope is that $V_{\mathcal{F}}$ can be still bounded up to some coverage condition that measures how $\pi_{e}$ deviates from $\pi_{b}$.
Unfortunately, generalizing the above equation to the off-policy case raises issues: consider the solution
$\textstyle V_{\mathcal{F}}(f_{h})=R(f_{h})\cdot\prod_{h^{\prime}=h}^{H}\frac{\pi_{e}(a_{h^{\prime}}\mid o_{h^{\prime}})}{\pi_{b}(a_{h^{\prime}}\mid o_{h^{\prime}})},$
whose correctness can be verified by the principle of importance sampling: for each $s_{h}$, $\left(M_{\mathcal{F},h}V_{\mathcal{F}}\right)(s_{h})=$
$~{}\textstyle\sum_{f_{h}}R(f_{h})\operatorname{\Pr}_{\pi_{b}}(f_{h}\mid s_{h})\cdot\prod_{h^{\prime}=h}^{H}\frac{\pi_{e}(a_{h^{\prime}}\mid o_{h^{\prime}})}{\pi_{b}(a_{h^{\prime}}\mid o_{h^{\prime}})}$
$=$ $~{}\textstyle\sum_{f_{h}}R(f_{h})\operatorname{\Pr}_{\pi_{e}}(f_{h}\mid s_{h})=V_{{\mathcal{S}}}^{\pi_{e}}(s_{h}).$
Although we obtain a valid solution, the construction involves cumulative action importance weights, which erases the superiority of the framework over IS as discussed in the introduction.
Two additional remarks:
  1. 1.
    While our intuition tells that $V_{\mathcal{F}}$ may be bounded with additional coverage assumptions, standard assumptions such as latent state coverage are clearly irrelevant here: latent state coverage compares $d^{\pi_{e}}(s_{h})$ and $d^{\pi_{b}}(s_{h})$, which are occupancy distributions induced by the initial latent distribution $d_{1}$ and the dynamics between time $1$ and $h$. In contrast, when we use importance weights in the above construction, it is to correct the distribution mismatch after step $h$ and conditioned on $s_{h}$; the $s_{h}$ conditioning completely blocks the influence of the dynamics between time $1$ and $h$. Therefore, what we need is a novel notion of coverage that is not seen in MDPs, which concerns the future rather than the past, and we will propose such a concept shortly.
  2. 2.
    As a side product, the construction shows that Eq.(6) always has a solution (as long as Assumption 1 holds), even if $M_{\mathcal{F},h}$ does not have full row-rank.
@@
Another direction is to simply treat Eq.(6) as a linear system. Given that the system is under-determined, we can use pseudo-inverse: [^4] All covariance-like matrices in the paper, such as $M_{\mathcal{F},h}M_{\mathcal{F},h}^{\top}$ here, are invertible under Assumption 2.
$V_{\mathcal{F},h}=M_{\mathcal{F},h}^{\top}\left(M_{\mathcal{F},h}M_{\mathcal{F},h}^{\top}\right)^{-1}V_{{\mathcal{S}},h}^{\pi_{e}}.$
In this case, $\|V_{\mathcal{F}}\|_{\infty}$ can be bounded using $1/\sigma_{\min}(M_{\mathcal{F},h})$, where $\sigma_{\min}$ denotes the smallest singular value. In fact, closely related quantities have appeared in the recent POMDP literature; for example, in the online setting, Liu et al. (2022) used $\sigma_{\min}$ of an action-conditioned variant of $M_{\mathcal{F},h}$ as a complexity parameter for online exploration in POMDPs. Unfortunately, these quantities suffer from scaling issues: $1/\sigma_{\min}(M_{\mathcal{F},h})$ is guaranteed to be misbehaved if the process is sufficiently stochastic.
@@
[…]
In this example, $C_{\textrm{stoch}}$ measures how the distribution of $f_{h}$ under $\pi_{b}$ deviates multiplicatively from a uniform distribution over $\mathcal{F}_{h}$, and an even moderately stochastic $\pi_{b}$ and emission process $\mathbb{O}$ will lead to small $C_{\textrm{stoch}}$ (or an exponential one that grows slower than $(OA)^{H-h+1}$), which implies an exponentially large $1/\sigma_{\min}(M_{\mathcal{F},h})$.
Minimum Weighted $2$-Norm Solution.
@@
$V_{\mathcal{F},h}=Z_{h}^{-1}M_{\mathcal{F},h}^{\top}\Sigma_{\mathcal{F},h}^{-1}V_{{\mathcal{S}},h}^{\pi_{e}},$ (7)
$\textrm{where~{}}\Sigma_{\mathcal{F},h}:=M_{\mathcal{F},h}Z_{h}^{-1}M_{\mathcal{F},h}^{\top}.$
$\Sigma_{\mathcal{F},h}\in\mathbb{R}^{S\times S}$ plays an important role in this construction. Recall that its counterpart in the pseudo-inverse solution, namely $M_{\mathcal{F},h}M_{\mathcal{F},h}^{\top}$, has scaling issues (Example 1), that even its largest eigenvalue can decay exponentially with $H-h+1$. In contrast, $\Sigma_{\mathcal{F},h}$ is very well-behaved in its magnitude, as shown below. Furthermore, while $M_{\mathcal{F},h}^{\top}$ on the left is now multiplied by $Z_{h}^{-1}$ which can be exponentially large, $Z_{h}^{-1}M_{\mathcal{F},h}^{\top}$ together is still well-behaved; see proof in Appendix C.2.
@@
Suppose $f_{h}$ always reveals $s_{h}$, in the sense that for any $j\in\mathcal{F}_{h}$, $\Pr_{\pi_{b}}(f_{h}=j\mid s_{h}=i)$ is only non-zero for a single $i\in[S]$, and zero for all other latent states. Then, $\Sigma_{\mathcal{F},h}=\mathbf{I}$, the identity matrix. Furthermore, $V_{\mathcal{F}}$ from Eq.(7) satisfies $\|V_{\mathcal{F}}\|_{\infty}\leq H.$ See Appendix C.3 for details.
The example shows an ideal case where $\sigma_{\min}(\Sigma_{\mathcal{F},h})=\sigma_{\max}(\Sigma_{\mathcal{F},h})=1$, when the future is “revealing” for step $h$ [^5] More generally, $\Sigma_{\mathcal{F},h}$ is the confusion matrix when we predict $s_{h}$ from $f_{h}$ using the posterior based on a uniform prior over ${\mathcal{S}}_{h}$. This can happen when e.g., the last observation $o_{H}$ reveals the identity of an earlier latent state $s_{h}$. Note that in this case, $M_{\mathcal{F},h}M_{\mathcal{F},h}^{\top}$ can still have poor scaling, if the actions and observations between step $h$ and $H$ are sufficiently stochastic, which shows how the weighted 2-norm solution and its analysis improve over the pseudo-inverse one. More generally, $\sigma_{\min}(\Sigma_{\mathcal{F},h})$ serves as a measure of how the distribution of future $f_{h}$ helps reveal the latent state $s_{h}$, an intuition also assigned to $\sigma_{\min}(M_{\mathcal{F},h})$ in prior works; $\Sigma_{\mathcal{F},h}$ just enjoys significantly better properties in comparison.
We now break down the boundedness of Eq.(7) into more interpretable assumptions in a way standard in the linear MDP literature (Jin et al., 2020). As we will see later, while such a breakdown gives reasonable guarantees when $S$ is small, it does not best fit the structure of POMDPs and can be further improved (Section 5.2). Nevertheless, we still present them below to help develop intuitions for later improvements and connect to the MDP literature.
@@
Assume the following holds for any $f_{h}$: $\|\mathbf{u}(f_{h})/Z(f_{h})\|^{2}_{\Sigma_{\mathcal{F},h}^{-1}}\leq C_{\mathcal{F},U}$.
@@
See proof in Appendix C.4. The proposition follows directly by $V_{\mathcal{F}}(f_{h})=(\frac{\mathbf{u}(f_{h})^{\top}}{Z(f_{h})}\Sigma_{\mathcal{F},h}^{-1/2})(\Sigma_{\mathcal{F},h}^{-1/2}V_{{\mathcal{S}},h}^{\pi_{e}})$ and Hölder’s. We now provide interpretations on the assumptions:
Interpretation of $L_{2}$ Outcome Coverage.
Assumption 7 requires that the weighted covariance matrix $\Sigma_{\mathcal{F},h}$ covers the direction of $V_{{\mathcal{S}},h}^{\pi_{e}}$ well. As a sanity check, it is always bounded in the on-policy case:
@@
(We will address the $S$ factor later; for now we treat $S$ as a small quantity.) When $1/\sigma_{\min}(\Sigma_{\mathcal{F},h})$ is bounded, $\|V_{{\mathcal{S}},h}^{\pi_{e}}\|_{\Sigma_{\mathcal{F},h}^{-1}}$ is immediately bounded. Even when $\sigma_{\min}(\Sigma_{\mathcal{F},h})$ is small, $C_{\mathcal{F},2}$ can still be small as long as $V_{{\mathcal{S}}}^{\pi_{e}}$ does not activate those small eigenvalue directions.
Notably, mathematically similar coverage assumptions are also found in the linear MDP literature. For example, with state-action feature $\phi_{h}$ for time step $h$, a very tight coverage parameter for linear MDPs (Yin et al., 2022) is
$\mathbb{E}_{\pi_{e}}[\phi_{h}]^{\top}\mathbb{E}_{\pi_{b}}[\phi_{h}\phi_{h}^{\top}]^{-1}\mathbb{E}_{\pi_{e}}[\phi_{h}].$
Despite the mathematically similarity, there are important high-level differences between these notions of coverage:
@@
    As mentioned earlier, MDP coverage is concerned with the dynamics before $h$, whereas our outcome coverage concerns that after $H$. Relatedly, MDP coverage depends on the initial distribution (which our outcome coverage does not depend on), and our coverage depends on the reward function through $V_{\mathcal{F}}$ (which MDP coverage does not explicitly depend on). In Section 5, we will discuss our other coverage assumption (belief coverage), which is more similar to the MDP coverage in that they are both concerned with the past.
@@
    The linear MDP coverage assumption is a refinement of state-density ratio using the knowledge of the function class (Chen & Jiang, 2019; Song et al., 2022), which gives its linear-algebraic structure. [^6] While the MDP-reduction approach discussed in Section 1 can also benefit from the linear structure of function classes, this requires encoding entire history into low-dimensional linear features, and is difficult without strong prior knowledge. In comparison, the linear structure of our outcome-coverage assumption comes directly from the internal structure of POMDPs.
Interpretation of $\Sigma_{\mathcal{F},h}$ Regularity.
With Assumption 7 alone, we can already shown the boundedness of $\|V_{\mathcal{F},h}\|_{Z_{h}}$, which is a weighted 2-norm that can be useful in Bernstein-type concentration inequalities. However, to obtain $L_{\infty}$ bound we need an additional regularity condition on $\Sigma_{\mathcal{F},h}$. To interpret Assumption 8, let $x=\mathbf{u}(f_{h})/Z(f_{h})\in\Delta({\mathcal{S}})$, and assign a distribution over $x$ by sampling $f_{h}\sim\textrm{diag}(Z_{h}/S)$, then Assumption 8 becomes $x^{\top}\mathbb{E}_{Z_{h}/S}[xx^{\top}]^{-1}x\leq C_{\mathcal{F},U}S$ for any $x$ with non-zero probability. This is a common regularity assumption (Duan et al., 2020; Perdomo et al., 2023), requiring that no $x$ points to a direction in $\mathbb{R}^{S}$ alone without being joined by others from the distribution. While $1/\sigma_{\min}(\Sigma_{\mathcal{F},h})$ can bound this quantity, the other way around is not true, implying that Assumption 8 is weaker.

### v2

@@ 3 Future-dependent Value Functions
$|J(\pi_{e})-$ $\mathbb{E}_{\mathcal{D}}[\widehat{V}(f_{1})]|\leq cH\max\{C_{\mathcal{V}}+1,C_{\Xi}\}\cdot\mathrm{IV}(\mathcal{V})\mathrm{Dr}_{\mathcal{V}}[d^{\pi_{e}},d^{\pi_{b}}]\sqrt{\frac{C_{\mu}\log\frac{|\mathcal{V}||\Xi|}{\delta}}{n}},$
where $c$ is an absolute constant, [^3] The value of $c$ can differ in each occurrence, and we reserve the symbol $c$ for such absolute constants. and $C_{\mathcal{V}}:=\max_{V\in\mathcal{V}}\|V\|_{\infty}$, $C_{\Xi}:=\max_{\xi\in\Xi}\|\xi\|_{\infty}$,
$\mathrm{IV}(\mathcal{V})$ $:=\max_{h}\sup_{V\in\mathcal{V}}\sqrt{\frac{\mathbb{E}_{\pi_{b}}\left[\left(\mathcal{B}^{{\mathcal{S}}}V\right)(s_{h})^{2}\right]}{\mathbb{E}_{\pi_{b}}\left[\left(\mathcal{B}^{\mathcal{H}}V\right)(\tau_{h})^{2}\right]}},\quad\mathrm{Dr}_{\mathcal{V}}[d^{\pi_{e}},d^{\pi_{b}}]:=\max_{h}\sup_{V\in\mathcal{V}}\sqrt{\frac{\mathbb{E}_{\pi_{e}}\left[\left(\mathcal{B}^{{\mathcal{S}}}V\right)(s_{h})^{2}\right]}{\mathbb{E}_{\pi_{b}}\left[\left(\mathcal{B}^{{\mathcal{S}}}V\right)(s_{h})^{2}\right]}}.$
See proof in Appendix C.2. Among the objects that appear in the bound, some are mundane and expected (e.g., horizon $H$, the complexities of function classes $\log|\mathcal{V}||\Xi|$, etc.). We now focus on the important ones, which reveals the open questions we shall investigate next:
@@
    $\mathrm{Dr}_{\mathcal{V}}[d^{\pi_{e}},d^{\pi_{b}}]$ measures the coverage of $\pi_{b}$ over $\pi_{e}$ on the latent state space ${\mathcal{S}}$, as the expression inside square-root can be bounded by $\max_{s_{h}}d^{\pi_{e}}(s_{h})/d^{\pi_{b}}(s_{h})$.
@@
    (Q2) $\mathrm{IV}(\mathcal{V})$ appears because we use $\mathcal{B}^{\mathcal{H}}V$ as a proxy for $\mathcal{B}^{{\mathcal{S}}}V$, which is only a linear measure of the latter, $(\mathcal{B}^{\mathcal{H}}V)(\tau_{h})=\langle\mathbf{b}(\tau_{h}),\mathcal{B}^{{\mathcal{S}}}_{h}V\rangle$. $\mathrm{IV}(\mathcal{V})$ reflects the conversion ratio between them. Uehara et al. (2022a) pointed out that the value is finite if $M_{\mathcal{H},h}$ has full-row rank ($S$), but a quantitative understanding is missing.
The rest of the paper proposes new coverage assumptions, analyses, and algorithms to answer the above questions, providing deeper understanding on the mathematical structure of OPE in POMDPs.
@@ 4 Boundedness of FDVFs
We start with Q1: when are FDVFs bounded? Recall from Definition 3 that a FDVF at level $h$, $V_{\mathcal{F},h}$, is any function satisfying $M_{\mathcal{F},h}\times V_{\mathcal{F},h}=V_{{\mathcal{S}},h}^{\pi_{e}}$, where $M_{\mathcal{F},h}$ is the outcome matrix with $(M_{\mathcal{F},h})_{i,j}=\operatorname{\Pr}_{\pi_{b}}(f_{h}=j\mid s_{h}=i)$. Since the equation can have many solutions and we only need to find one of them, it suffices to provide an explicit construction of a $V_{\mathcal{F},h}$ and show it is well bounded. Also note that the equations for different $h$ are independent of each other, so we can construct $V_{\mathcal{F},h}$ for each $h$ separately.
We first describe two natural constructions, both of which yield exponentially large upper bounds.
@@
Uehara et al. (2022a) provided a cruel bound on $\|V_{\mathcal{F}}\|_{\infty}$ which scales with $1/\sigma_{\min}(M_{\mathcal{F},h})$, which we show shortly is an exponential-in-length quantity. However, it is not hard to notice that in certain benign cases, $V_{\mathcal{F}}$ does not have to blow-up exponentially and has obviously bounded constructions.
As a warm-up, consider the on-policy case of $\pi_{b}=\pi_{e}$, where a natural solution is $V_{\mathcal{F}}(f_{h})=R^{+}(f_{h}):=\sum_{h^{\prime}=h}^{H}R(o_{h^{\prime}},a_{h^{\prime}}).$ Here $R^{+}(f_{h})$ simply adds up the Monte-Carlo rewards in future $f_{h}$, which is bounded in $[0,H]$. The construction trivially satisfies $\mathbb{E}_{\pi_{b}}\left[V_{\mathcal{F}}(f_{h})\mid s_{h}\right]=V_{{\mathcal{S}}}^{\pi_{e}}(s_{h})$ when $\pi_{b}=\pi_{e}$. In the more general setting of $\pi_{b}\neq\pi_{e}$, the hope is that $V_{\mathcal{F}}$ can be still bounded up to some coverage condition that measures how $\pi_{e}$ deviates from $\pi_{b}$.
Unfortunately, generalizing the above equation to the off-policy case raises issues: consider the solution $V_{\mathcal{F}}(f_{h})=R^{+}(f_{h})\cdot\prod_{h^{\prime}=h}^{H}\frac{\pi_{e}(a_{h^{\prime}}\mid o_{h^{\prime}})}{\pi_{b}(a_{h^{\prime}}\mid o_{h^{\prime}})},$ whose correctness can be verified by importance sampling. Despite its validity, the construction involves cumulative action importance weights, which erases the superiority of the framework over IS as discussed in the introduction.
@@
Another direction is to simply treat Eq.(1) as a linear system. Given that the system is under-determined, we can use pseudo-inverse: [^4] All covariance-like matrices in the paper, such as $M_{\mathcal{F},h}M_{\mathcal{F},h}^{\top}$ here, are invertible under Assumption 2. $V_{\mathcal{F},h}=M_{\mathcal{F},h}^{\top}\left(M_{\mathcal{F},h}M_{\mathcal{F},h}^{\top}\right)^{-1}V_{{\mathcal{S}},h}^{\pi_{e}}.$ In this case, $\|V_{\mathcal{F}}\|_{\infty}$ can be bounded using $1/\sigma_{\min}(M_{\mathcal{F},h})$, where $\sigma_{\min}$ denotes the smallest singular value. In fact, closely related quantities have appeared in the recent POMDP literature; for example, in the online setting, Liu et al. (2022a) used $\sigma_{\min}$ of an action-conditioned variant of $M_{\mathcal{F},h}$ as a complexity parameter for online exploration in POMDPs. Unfortunately, these quantities suffer from scaling issues: $1/\sigma_{\min}(M_{\mathcal{F},h})$ is guaranteed to be misbehaved if the process is sufficiently stochastic.
@@
If $\Pr_{\pi_{b}}(f_{h}|s_{h})\leq\frac{C_{\textrm{stoch}}}{(OA)^{H-h+1}}$, $\forall f_{h},s_{h}$, then $\sigma_{\min}(M_{\mathcal{F},h})\leq C_{\textrm{stoch}}\sqrt{S}/(OA)^{\frac{H-h+1}{2}}$.
In this example, $C_{\textrm{stoch}}$ measures how the distribution of $f_{h}$ under $\pi_{b}$ deviates multiplicatively from a uniform distribution over $\mathcal{F}_{h}$, and an even moderately stochastic $\pi_{b}$ and emission process $\mathbb{O}$ will lead to small $C_{\textrm{stoch}}$, which implies an exponentially large $1/\sigma_{\min}(M_{\mathcal{F},h})$.
4.1 Minimum Weighted 2-Norm Solution and $L_{2}$ Outcome Coverage
@@
$V_{\mathcal{F},h}=Z_{h}^{-1}M_{\mathcal{F},h}^{\top}\Sigma_{\mathcal{F},h}^{-1}V_{{\mathcal{S}},h}^{\pi_{e}},~{}~{}~{}\textrm{where~{}}\Sigma_{\mathcal{F},h}:=M_{\mathcal{F},h}Z_{h}^{-1}M_{\mathcal{F},h}^{\top}.$ (4)
$\Sigma_{\mathcal{F},h}\in\mathbb{R}^{S\times S}$ plays an important role in this construction. Recall that its counterpart in the pseudo-inverse solution, namely $M_{\mathcal{F},h}M_{\mathcal{F},h}^{\top}$, has scaling issues (Example 1), that even its largest eigenvalue can decay exponentially with $H-h+1$. In contrast, $\Sigma_{\mathcal{F},h}$ is very well-behaved in its magnitude, as shown below. Furthermore, while $M_{\mathcal{F},h}^{\top}$ on the left is now multiplied by $Z_{h}^{-1}$ which can be exponentially large, $Z_{h}^{-1}M_{\mathcal{F},h}^{\top}$ together is still well-behaved; see proof in Appendix D.2.
@@
Suppose $f_{h}$ always reveals $s_{h}$, in the sense that for any $j\in\mathcal{F}_{h}$, $\Pr_{\pi_{b}}(f_{h}=j\mid s_{h}=i)$ is only non-zero for a single $i\in[S]$, and zero for all other latent states. Then, $\Sigma_{\mathcal{F},h}=\mathbf{I}$, the identity matrix. Furthermore, $V_{\mathcal{F}}$ from Eq.(4) satisfies $\|V_{\mathcal{F}}\|_{\infty}\leq H.$ See Appendix D.3 for details.
The example shows an ideal case where $\sigma_{\min}(\Sigma_{\mathcal{F},h})=\sigma_{\max}(\Sigma_{\mathcal{F},h})=1$, when the future fully determines $s_{h}$. This can happen when the last observation $o_{H}$ reveals the identity of an earlier latent state $s_{h}$. Note that in this case, $M_{\mathcal{F},h}M_{\mathcal{F},h}^{\top}$ can still have poor scaling if the actions and observations between step $h$ and $H$ are sufficiently stochastic, which shows how the weighted 2-norm solution and analysis improve over the pseudo-inverse one. More generally, $\Sigma_{\mathcal{F},h}$ is the confusion matrix of making posterior predictions of $s_{h}$ from $f_{h}$ based on a uniform prior over ${\mathcal{S}}_{h}$ (see Appendix B.8 for how to incorporate different priors) with $\mathbf{u}(f_{h})^{\top}/Z(f_{h})$ being the posterior, and $\sigma_{\min}(\Sigma_{\mathcal{F},h})$ serves as a measure of how the distribution of future $f_{h}$ helps reveal the latent state $s_{h}$.
We now break down the boundedness of Eq.(4) into more interpretable assumptions.
@@
Assume for any $f_{h}$: $\|\mathbf{u}(f_{h})/Z(f_{h})\|^{2}_{\Sigma_{\mathcal{F},h}^{-1}}\leq C_{\mathcal{F},U}$.
@@
See proof in Appendix D.4. Assumption 7 requires that the weighted covariance matrix $\Sigma_{\mathcal{F},h}$ covers the direction of $V_{{\mathcal{S}},h}^{\pi_{e}}$ well. As a sanity check, it is always bounded in the on-policy case:
@@
Notably, mathematically similar coverage assumptions are also found in the linear MDP literature. For example, with state-action feature $\phi_{h}$ for time step $h$, a very tight coverage parameter for linear MDPs is $\|\mathbb{E}_{\pi_{e}}[\phi_{h}]\|_{\mathbb{E}_{\pi_{b}}[\phi_{h}\phi_{h}^{\top}]^{-1}}^{2}$ (Zanette et al., 2021). Despite the mathematically similarity, there are important high-level differences between these notions of coverage:
@@
    As mentioned earlier, MDP coverage is concerned with the dynamics before step $h$, whereas our outcome coverage concerns that after $h$. Relatedly, MDP coverage depends on the initial distribution (which our outcome coverage does not depend on), and our coverage depends on the reward function through $V_{\mathcal{F}}$ (which MDP coverage does not explicitly depend on). In Section 5, we will discuss our other coverage assumption (belief coverage), which is more similar to the MDP coverage in that they are both concerned with the past.
@@
    The linear MDP coverage assumption is a refinement of state-density ratio using the knowledge of the function class (Chen and Jiang, 2019; Song et al., 2022). In comparison, the linear structure of our outcome-coverage assumption comes directly from the internal structure of POMDPs.
4.2 Addressing $S$ dependence via $L_{1}/L_{\infty}$ Hölder and $L_{\infty}$ Outcome Coverage
Example 3 shows that even in the on-policy case, $C_{\mathcal{F},V}$ may depend on $S$ which makes the assumption only meaningful for finite and small ${\mathcal{S}}$. In fact, we showed earlier that $V_{\mathcal{F}}=R^{+}$ is a natural and obvious $L_{\infty}$-bounded solution for $\pi_{b}=\pi_{e}$, but this is not recovered by the construction in Eq. (4). We also need an additional regularity Assumption 8.
As it turns out, these undesired properties arise because $L_{2}$ Hölder—which is natural for linear MDP settings mentioned above—fails to leverage the $L_{1}$ normalization of $\mathbf{u}(f_{h})^{\top}/Z(f_{h})$ (Proposition 4, Claim 2) and is loose for POMDPs; see Appendix B.5 for further details. A better choice is $L_{1}/L_{\infty}$ Hölder, motivating the $L_{\infty}$ coverage assumption below, which requires a slightly different construction of $V_{\mathcal{F}}$. These definitions may seem mysterious or even counterintuitive; it will be easier to explain the intuitions when we get to their counterparts for belief coverage in Section 5.2.
Construction of $V_{\mathcal{F}}$
Define $Z^{R}(f_{h}):=Z(f_{h})/R^{+}(f_{h})$, and we use $Z^{R}$ to replace $Z$ in Eq.(4):
$V_{\mathcal{F},h}=(Z_{h}^{R})^{-1}M_{\mathcal{F},h}^{\top}(\Sigma_{\mathcal{F},h}^{R})^{-1}V_{{\mathcal{S}},h}^{\pi_{e}},~{}~{}~{}\textrm{where~{}}\Sigma_{\mathcal{F},h}^{R}:=M_{\mathcal{F},h}(Z_{h}^{R})^{-1}M_{\mathcal{F},h}^{\top}.$ (5)
Assumption 9 ($L_{\infty}$ outcome coverage).
Assume for all $h$, $\|(\Sigma_{\mathcal{F},h}^{R})^{-1}V_{{\mathcal{S}}}^{\pi_{e}}\|_{\infty}\leq C_{\mathcal{F},\infty}.$
Lemma 5.
Under 9, $\|V_{\mathcal{F}}\|_{\infty}\leq HC_{\mathcal{F},\infty}$. See proof in Appendix E.7.
In Appendix B.5 we show that the construction shares similar properties to Eq.(4) in the scenario of Example 2. On the other hand, it has better scaling properties w.r.t. $S$ and does not additionally require a regularity assumption like Assumption 8. In the on-policy case, Eq.(5) exactly recovers $V_{\mathcal{F}}=R^{+}$, a property that Eq.(4) does not enjoy; see Appendix E.8 for details.
Example 4.
When $\pi_{e}=\pi_{b}$, $(\Sigma_{\mathcal{F},h}^{R})^{-1}V_{{\mathcal{S}}}^{\pi_{e}}=\mathbf{1}$, thus 9 holds with $C_{\mathcal{F},\infty}=1$ (c.f. $C_{\mathcal{F},V}\leq H\sqrt{S}$ in Example 3). Furthermore, the construction in Eq.(5) is exactly $V_{\mathcal{F}}=R^{+}$.